Fach

\underline{Fach ID}  Stundendotation Spezialzimmer
Deutsch 4 —
Französisch 4  —
Englisch  4  —
Mathe  4  —
Biologie  2  Praktikumszimmer B
Physik  2  Praktikumszimmer P
Chemie  2  Chemielabor
Geographie 2  —
Geschichte 2  —
Informatik  2  —
WR  4  —
PPP  4  —
BG  2  Zeichensaal
Musik  2  —
Sport  3  Turnhalle